Circular No. 7131

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1999az, 1999ba, 1999bb L. G. Strolger, University of Michigan; and R. C. Smith, Cerro Tololo

Interamerican Observatory (CTIO), report for the Nearby Galaxies Supernova Search team (cf. IAUC 7125) and the Supernova Cosmology Project (cf. IAUC 7125, 7128) the discovery of three supernovae in R-band images taken with the Kitt Peak 0.9-m telescope (+ MOSAIC camera):

SN 1999 UT α_{2000} δ_{2000} R Offset
1999az Feb. 23 10^{h}18^{m}56.^{s}37 −10^{o}25′18′′.1 20.5 near center
1999ba Feb. 23 10 19 19.37 − 2 49 39.0 20.0 near center
1999bb Feb. 23 11 05 16.47 − 6 01 32.5 20.6  ?

Based on spectra taken by Strolger and Smith on Mar. 17 UT at the CTIO 4-m telescope, P. Nugent (Lawrence Berkeley Laboratory) and M. Phillips (Las Campanas Observatory) identify SN 1999az as a type-Ia supernova (redshift ∼ 0.15), observed ∼ 1 month past maximum brightness, and SN 1999ba as a type-Ia event (redshift ∼ 0.09), observed between 20 and 30 days past maximum brightness. Both SNe 1999az and 1999ba are slightly offset from the core of their apparent host galaxies. Based on spectra taken by Strolger and Smith on Mar. 18 at the same telescope, SN 1999bb has been identified as a type-II supernova, between 20 and 30 days past maximum brightness (redshift ∼ 0.18); no host is apparent in the discovery images. Finding charts for the three objects (identified as nb-125, nb-092, and nb-088, respectively) can be obtained

from http://panisse.lbl.gov/nearsearch/fcharts/index.html. The

Nearby Galaxies Supernova Search team was also involved in the discoveries of SNe 1999au, 1999av, and 1999aw reported on IAUC 7130.

S4 0954+65 C. M. Raiteri, M. Villata, G. Massone, G. De Francesco, L. Lanteri, L.

Ostorero, M. Puccio, and G. Sobrito, Osservatorio Astronomico di Torino; and G. Tosti, Università di Perugia, report: “S4 0954+65 (cf. Pauliny-Toth et al. 1978, A.J. 83, 451) is undergoing a major optical outburst (see Blazar

Data News 14 and 19, http://bldata.pg.infn.it/news index.html);

CCD data taken with the Torino 1.05-m and the Perugia 0.4-m telescopes on Mar. 24.97 and 25.00 UT show a new historical maximum at R = 14.63 ± 0.03. Observations at other wavelengths are strongly encouraged.”

1999 March 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green